OpenStudy (anonymous):

help anyone??? Show the work to solve |x+3| +7 less than or equal to 10

OpenStudy (turingtest):

\[\left| x+3 \right|+7 \le 10\] \[\left| x+3 \right| \le 3\]\[-3 \le x+3 \le 3\]\[0 \le x \le 6\]

OpenStudy (turingtest):

sorry the last part should be\[-6 \le x \le 0\]

OpenStudy (anonymous):

what hapened to the 7?

OpenStudy (turingtest):

at the beginning while we still had absolute value signs I subtracted it.

OpenStudy (turingtest):

that's why from there on we have < or = to 3

OpenStudy (anonymous):

oh okay

OpenStudy (turingtest):

|x+3|+7≤10 subtract 7 from both sides: |x+3|≤3 −3≤x+3≤3 −6≤x≤0 is how I was supposed to write it